[ If one car costs $14,300 and it gets 35 mpg. And the other car costs [$15,400] and gets 40 mpg. Assuming that gas costs $3.75 per gallon [at what number of driven miles] would the two cars [be of equal value?] ]
Answer by TimothyLamb(4379)

3.75 $/gallon
---
x = miles driven
---
car a:
m = 3.75 $/gal * (1/35) gal/mile = 0.1071428 $/mile (fuel cost per mile)
---
car b:
m = 3.75 $/gal * (1/40) gal/mile = 0.09375 $/mile (fuel cost per mile)
---
cost of ownership per mile:
---
a(x) = 0.1071428x + 14300
b(x) = 0.09375x + 15400
---
break even point:
0.1071428x + 14300 = 0.09375x + 15400
0.1071428x - 0.09375x = 15400 - 14300
0.0133928x = 1100
---
x = 82134 miles
---
Answer:
the more fuel efficient car-a, which is $1100 more expensive to purchase, will not match the less fuel efficient car-b in terms of operating cost per mile, until the cars have each been driven 82,134 miles. In other words, the more fuel efficient car-a will cost more to operate per mile than car-b, until both cars have each been driven 82,134 miles, at which time car-a will cost less to operate per mile than car-b.
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php